Bigbie Corp. issued a three-year bond a year ago with a coupon of 8 percent. The bond pays interest semiannually. If the yield to maturity on this bond is 7.4 percent, what is the price of the bond? Round your answer to 2 decimal places.
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 2 * 2 = 4 semi-annual periods, C = 8% *$1000/2 = $40 (semi-annually), i = 7.4%/2 = 3.7% (semi-annually)

1- (1+0.037) P = 40 * - 1000 (1 + 0.037) 0.037

P = $146.23 + $864.74

P = $1,010.97